Simplify and write your answer in scientific notation,show all your work (no calculator). (7.3*10^5)(2.0*10^-2)
Answer by stanbon(75887)

Simplify and write your answer in scientific notation,show all your work (no calculator). (7.3*10^5)(2.0*10^-2)
-----
= 7.3*2 *10^3
-----
= 14.6 * 10^3
---
= 14.6/10 * 10*10^3
-----
= 1.46 + 10^4
===================
Cheers,
Stan H.
===================